# How Cohesive Are Community Search Results on Online Social Networks?: An Experimental Evaluation

### v2

@@ 7. Experimental Evaluation
In this section, we conduct an in-depth evaluation of the representative CS algorithms. We first assess the structural cohesiveness of their identified communities, followed by a detailed evaluation of their quality based on five psychology-informed cohesiveness measures. Learning-based algorithms were run on a Linux server with an Intel Xeon Platinum 8255C CPU (2.50GHz) and an RTX 3080 GPU (10GB memory). The remaining experiments were conducted on a Linux server with an Intel Xeon(R) Platinum 8352V CPU (2.10GHz) and 120GB memory. Recall that we do not focus on efficiency and scalability analysis here, as our goal is to understand the effectiveness of CS algorithms. The reader may refer to (Fang et al., 2020) for studies related to the former.
@@ 7.1. Structural Cohesiveness
Table 6 reports the results. We can make the following observations: (1) Although ALS, Repeeling+, and TransZero-LS have high $Q_{hit}$, they tend to generate larger yet sparser communities, characterized by larger $\mathbf{d}$, larger sizes, and lower $Deg_{min}$; (2) I2ACSM identifies very small communities across all datasets; (3) Compared to aforementioned algorithms, WCF-CRC, CSD, and ST-Exa are better at identifying structurally cohesive communities. However, WCF-CRC and CSD generally achieve lower hit rates, indicating difficulty in identifying target communities. Among all algorithms, ST-Exa successfully returns communities for all queries and demonstrates superior performance across all datasets, partially due to its size-constrained nature. Note that the communities (or node lists) returned by TransZero-GS are disconnected, likely due to insufficient node features during search and ignoring edge connections when selecting candidates. Hence, its results are left blank.
The poor performance of ALS may result from the smaller time scales adopted in our datasets, which hinder its ability to find nodes with high temporal proximity. However, the original paper does not discuss the time scale and its impact on CS results. Additionally, the identification of extremely large or small communities may be due to graph sparsity after multigraph transformations. The unpromising performance of the two learning-based methods may stem from the insufficient node features in our datasets, limiting CSGphormer’s ability to learn adequate features for calculating community scores. On the other hand, this highlights these methods’ stringent data requirements and their limitations when applied to datasets not specifically designed for model training.
@@ 7.2. Psychology-informed Cohesiveness
Enjoyment Index (EI). The top row of Figure 4 illustrates that community cohesiveness, as indicated by EI values, differs depending on the algorithm used and is affected by the importance assigned to recent activities. For example, communities identified by Repeeling+ on the BTW show higher enjoyment when $\lambda\in\{0.0001,0.0005,0.001\}$. However, increasing the weight of recent activities reduces enjoyment due to recent negative interactions. Notably, communities identified by CSD mostly exhibit positive EI values across all datasets and are less affected by the $\lambda$.
@@
We also examine the impact of a different time-decay function on EI, as shown in the bottom row of Figure 4. The results use a polynomial decay function, $\Phi(t-t_{cur})=(t_{cur}-t+1)^{-\mu}$ ($\mu>0$), which models a slower decline in the importance of past activities (Cormode et al., 2009). We vary $\mu$ from 0.5 to 2 in steps of 0.5. With a stronger and longer-lasting impact of recent interactions, EI values become more consistent across decay rates, and communities identified by CSD still largely maintain positive EI scores. Overall, none of the algorithms retrieved communities with consistently high EI scores.
@@
Sentimental Interaction Tendency (SIT). SIT scores are sensitive to decay rates and the choice of decay functions. Nevertheless, the SIT results are distinguishably different from the EI results. Figure 5 shows results on all datasets. On BTW, as $\lambda$ increases, no algorithms effectively capture mutual relationships among users. Furthermore, results on C144 reveal that communities with predominantly positive EI can still include negative mutual relationships (negative SIT), and vice versa. Although CSD and Repeeling+ consider directed edges, and ALS and WCF-CRC incorporate temporal information, their communities are still inadequate to achieve higher SIT scores.
@@
Comparative Enjoyment Degree (CED). Figure 6 plots the CED results. When $\lambda=0.0001$, some communities exhibit strong internal enjoyment, while others prefer external interactions, as indicated by the polarity of the CED scores. However, as the decay rate increases, the CED values for all communities converge towards zero, meaning the algorithms cannot find communities that increasingly prefer internal interactions over time. Similar to EI and SIT, CED is highly sensitive to decay function selection, which reflects the complexity of user interactions, as CED is calculated based on sentiments, timing, and where interactions occur. No algorithms consistently capture this complex interaction dynamics.
@@
Group Interaction Preference (GIP). The left plot in Figure 7 reports the GIP values. ALS, WCF-CRC, and TransZero-LS tend to identify communities that are highly interactive. However, as noted in Table 6, these high GIP values may result from the large community sizes returned by these algorithms, which include most interactions. Similar patterns are observed with higher GIP values for CSD communities on the C144 and C26 datasets and lower scores for I2ACSM communities across all datasets.
@@ 7.3. Impact of Sentiment Analysis Techniques
To examine the impact of sentiment analysis techniques on evaluation and to demonstrate the broad applicability of our cohesiveness metrics, we adopt a popular lexicon-based approach, VADER (Hutto and Gilbert, 2014), to relabel the datasets. We recompute sentiment-related cohesiveness measures (i.e., EI, SIT, and CED), and compare these results with those from Llama3-labeled datasets. Figure 8 presents two results on BTW datasets. In most cases, the cohesiveness scores are largely unaffected by sentiment analysis techniques. However, when the labeling technique changes to VADER, the ST-Exa and I2ACSM communities appear more sensitive, with their values either switching polarity (e.g., ST-Exa’s EI value on BTW) or showing obvious changes (e.g., ST-Exa’s and I2ACSM’s CED values on C144).
@@ 7.5. Case Study
Finally, we present a case study using the CC dataset to compare CS results from various algorithms, evaluating cohesiveness with our proposed measures and two additional structural metrics: $k$-core and $k$-truss. The structural metrics are computed after converting the CC into an undirected graph. Here, we present the results for node “158”, as most algorithms return its communities. To simplify visualization, we only display algorithms that return communities of size 5-30. If algorithms produce multiple communities under different parameters, the largest one is selected. Besides, self-loops within the networks are not displayed, and edge widths reflect the number of interactions between users (See Figure 10).
We observe that communities returned by different algorithms vary significantly. The first three algorithms identify larger communities with more internal connections. While the communities identified by I2ACSM and TransZero_LS exhibit frequent interactions between users, their overall connectivity is sparse. For structural cohesiveness, the first three communities are 2-core and 2-truss. However, despite being truss-based, Repeeling+ and I2ACSM fail to find more prominent $k$-truss communities. I2ACSM and TransZero_LS communities only meet the 1-core requirement. In other words, existing CS techniques fail to return consistent results for the same query node.
@@ 8. Reflections & Future Work
In this paper, we conduct the first study to extensively evaluate the cohesiveness of communities on online social networks produced by state-of-the-art CS algorithms. Our study revealed that communities identified by existing methods for the same query node lack consensus and their structural and psychological cohesiveness vary greatly, with no clear correlation between them. Despite the notion of cohesion being strongly rooted in social psychology theories, none of these generic techniques perform well on social networks w.r.t. the five psychological cohesiveness measures that are inspired by research in group cohesion. This is primarily because the existing one-size-fits-all cohesiveness measures of CS techniques do not systematically incorporate theories from social psychology in their design. This issue is further compounded by the fact that current CS techniques are either not assessed using “ground-truth” communities or rely on communities that are built without considering theories of group cohesion.
Despite these significant results, our evaluation has some limitations. Firstly, the design of our measures is constrained by sentiment analysis tools, and the simple three-class classification in our sentiment analysis may overlook nuanced sentiments inherent in specific comments. Secondly, our study is based on a single social network platform (i.e., X). However, given the fundamental limitations in the design of existing cohesiveness measures for human-interaction datasets (Section 4.2), we believe the key insights reported in this paper will also hold in other large online social networks. Thirdly, our study cannot use ground-truth communities for benchmarking because no public ground-truth datasets are created based on cohesion theories.
We believe the findings of our study can inspire a reevaluation of current community search algorithms by incorporating insights from social psychology, leading to the development of more effective solutions. Our proposed measures may be incorporated into search techniques to better identify cohesive communities in online social networks. In particular, it is paramount to create ground-truth community datasets for community search that are grounded in social psychology theories. This will not only enable the validation of our proposed psychology-informed cohesiveness measures but also assist existing CS techniques in developing more effective cohesiveness measures and algorithms for practical usage. Lastly, while this work concentrates on the community search problem, it is clear that our findings could also inspire a rethinking of various community-centric social search and analytics problems involving human interactions, such as community detection (Bedi and Sharma, 2016; Javed et al., 2018).

### v4

@@ 7. Experimental Evaluation
In this section, we conduct an in-depth evaluation of the representative CS algorithms and report key results. The reader may refer to (cohesiontech2025) for detailed results. We first assess the structural cohesiveness of their identified communities, then evaluate their quality using five psychology-informed cohesiveness measures. Learning-based algorithms were run on a Linux server with an Intel Xeon Platinum 8255C (2.50 GHz) and an RTX 3080 GPU (10 GB VRAM). Other experiments were conducted on a Linux server with an Intel Xeon Platinum 8352V (2.10 GHz) and 120 GB RAM. Recall that we do not focus on efficiency and scalability analysis here, as our goal is to understand the effectiveness of CS algorithms. The reader may refer to (Fang et al., 2020) for studies related to the former.
@@ 7.1. Structural Cohesiveness
Table 6 reports the results. We can make the following observations: (1) although ALS, Repeeling+, and TransZero-LS have high $Q_{hit}$, they tend to generate larger yet sparser communities characterized by greater degrees and sizes, and lower $Deg_{min}$ values; (2) I2ACSM identifies very small communities across all datasets; and (3) compared to aforementioned algorithms, WCF-CRC, CSD, and ST-Exa are better at identifying structurally cohesive communities. However, WCF-CRC and CSD generally achieve lower hit rates, indicating difficulty in identifying target communities. Among all algorithms, ST-Exa successfully returns communities for all queries and demonstrates superior performance across all datasets, partially due to its size-constrained nature. Note that the communities (or node lists) returned by TransZero-GS are disconnected, likely because its candidate selection ignores edge connections. Hence, its results are left blank.
The poor performance of ALS may stem from the smaller time scales used in our datasets, which hinder its ability to find nodes with greater temporal proximity. However, the original paper does not discuss the time scale and its impact on CS results. Additionally, identifying extremely large or small communities may arise from graph sparsity after multigraph transformations. The unpromising performance of the two learning-based methods may be attributable to insufficient node features in our datasets, limiting CSGphormer’s ability to learn adequate information for calculating community scores. On the other hand, this highlights these methods’ stringent data requirements and their limitations when applied to datasets not specifically designed for model training.
@@ 7.2. Psychology-informed Cohesiveness
Enjoyment Index (EI). The top row of Figure 4 illustrates that community cohesiveness, as indicated by EI values, varies across algorithms and is affected by the importance assigned to recent activities. For example, communities identified by Repeeling+ on the BTW show higher enjoyment when $\lambda\in\{0.0001,0.0005,0.001\}$. However, increasing the weight of recent activities reduces enjoyment due to recent negative interactions. Notably, communities identified by CSD mostly exhibit positive EI values across all datasets and are less affected by the $\lambda$.
@@
We also examine the impact of a different time-decay function on EI, as shown in the bottom row of Figure 4. Here, we use the polynomial decay function, $\Phi(t-t_{cur})=(t_{cur}-t+1)^{-\mu}$ ($\mu>0$), which models a slower decline in the importance of past activities (Cormode et al., 2009). We vary $\mu$ from 0.5 to 2 in steps of 0.5. With a stronger and more persistent impact of recent interactions, EI values become more consistent across decay rates, and communities identified by CSD still largely maintain positive EI scores. Overall, none of the algorithms retrieve communities with consistently high EI scores.
@@
Sentimental Interaction Tendency (SIT). SIT scores are sensitive to decay rates and the choice of decay functions. Nevertheless, the SIT results are distinguishably different from the EI results. Figure 5 shows results across the four datasets. On BTW, as $\lambda$ increases, no algorithms effectively capture mutual relationships among users. Furthermore, results on C144 reveal that communities with predominantly positive EI scores can still include negative mutual relationships (negative SIT values), and vice versa. Although CSD and Repeeling+ consider directed edges, and ALS and WCF-CRC incorporate temporal information, their communities are still inadequate to achieve higher SIT scores.
@@
Comparative Enjoyment Degree (CED). Figure 6 plots the CED results. When $\lambda=0.0001$, some communities exhibit strong internal enjoyment, while others prefer external interactions, as indicated by the polarity of the CED scores. However, as the decay rate increases, the CED values for all communities converge toward zero, meaning the algorithms cannot find communities that increasingly prefer internal interactions over time. Similar to EI and SIT, CED is highly sensitive to the choice of decay function, reflecting the complexity of user interactions, as CED is calculated based on sentiment, timing, and where interactions occur. No algorithms consistently capture these complex interaction dynamics.
@@
Group Interaction Preference (GIP). The left plot in Figure 7 reports the GIP values. ALS, WCF-CRC, and TransZero-LS tend to identify highly interactive communities. However, as noted in Table 6, these high GIP values may result from the large community sizes returned by these algorithms, which include most interactions. Similar patterns are observed with higher GIP values for CSD communities on the C144 and C26 datasets and lower scores for I2ACSM communities across all datasets.
@@ 7.3. Impact of Sentiment Analysis Techniques
To examine the impact of sentiment analysis techniques on evaluation and to demonstrate the broad applicability of our cohesiveness metrics, we adopt a popular lexicon-based approach, VADER (Hutto and Gilbert, 2014), to relabel the datasets. We recompute sentiment-related cohesiveness measures (i.e., EI, SIT, and CED), and compare these results with those from Llama3-labeled datasets. Figure 8 presents results across four datasets. In most cases, the cohesiveness scores are largely unaffected by sentiment analysis techniques. However, when the labeling technique changes to VADER, the ST-Exa and I2ACSM communities respond more strongly, with their values either switching polarity (e.g., ST-Exa’s EI value on BTW) or showing obvious changes (e.g., ST-Exa’s and I2ACSM’s CED values on C144).
@@ 7.5. Case Study
Finally, we present a case study using the CC dataset to compare CS results from various algorithms, evaluating cohesiveness with our proposed measures and two additional structural metrics: $k$-core and $k$-truss. The structural metrics are computed after converting the CC into an undirected graph. Here, we present the results for the query node “158”, as most algorithms return its communities. To simplify visualization, we only display algorithms that return communities of size 5-30. If algorithms produce multiple communities under different parameters, the largest one is selected. Besides, self-loops within the networks are not displayed, and edge widths reflect the number of interactions between users (See Figure 10).
We observe that communities returned by different algorithms vary significantly. The first three algorithms identify larger communities with more internal connections. While the communities from I2ACSM and TransZero-LS exhibit frequent interactions between users, their overall connectivity is sparse. For structural cohesiveness, the first three communities are 2-core and 2-truss. However, despite being truss-based, Repeeling+ and I2ACSM fail to find more prominent $k$-truss communities. I2ACSM and TransZero-LS communities only meet the 1-core requirement. In other words, existing CS techniques fail to return consistent results for the same query node.
@@ 8. Reflections & Future Work
In this paper, we conduct the first study to extensively evaluate the cohesiveness of communities in online social networks produced by state-of-the-art CS algorithms. Our study reveals that communities identified by existing methods for the same query node lack consensus, and their structural and psychological cohesiveness vary greatly, with no clear correlation between them. Despite the notion of cohesion being strongly rooted in social psychology theories, none of these generic techniques performs well on online social networks w.r.t. the five psychological cohesiveness measures that are inspired by research in group cohesion. This is primarily because the existing one-size-fits-all cohesiveness measures of CS techniques do not systematically incorporate theories from social psychology into their design. This issue is further compounded by the fact that current CS techniques are either not assessed using “ground-truth” communities or rely on communities that are built without considering theories of group cohesion.
Despite these significant results, our evaluation has some limitations. Firstly, the design of our measures is constrained by sentiment analysis tools, and the simple three-class classification in our sentiment analysis may overlook nuanced sentiments inherent in specific comments. Secondly, our study is based on a single social network platform (i.e., X). However, given the fundamental limitations in the design of existing cohesiveness measures for human-interaction datasets (Section 4.2), we believe that the key insights reported in this paper will also hold in other large online social networks. Thirdly, our study cannot use ground-truth communities for benchmarking because no public ground-truth datasets are created based on cohesion theories.
We believe that the findings of our study can inspire a reevaluation of current community search algorithms by incorporating insights from social psychology, leading to the development of more effective solutions. Our proposed measures may be incorporated into search techniques to better identify cohesive communities in online social networks. In particular, it is paramount to create ground-truth community datasets for community search that are grounded in social psychology theories. This will not only enable the validation of our proposed psychology-informed cohesiveness measures but also assist existing CS techniques in developing more effective cohesiveness measures and algorithms for practical usage. Lastly, while this work concentrates on the community search problem, it is clear that our findings could also inspire a rethinking of various community-centric social search and analytics problems involving human interactions, such as community detection (Bedi and Sharma, 2016; Javed et al., 2018).